Mr. Jack Tan, the Chief Executive Officer of RonSIM Industries Limited, is not clear on what the Company’s cost of capital is. RonSIM Industries Limited last dividend was $0.30 cents per share and the dividend is expected to grow at 5% per annum indefinitely. The stock is currently trading on Singapore Exchange at $4.70 per share.Case Background You are a financial analyst with Underwood & Company, a highly regarded management consultant company You have been seconded to RonSIM Industries Limited (the Company), a mainboard listed company, to undertake the evaluation of several proposals which are conceptualised to help the company new portfolio of investments. understand its financial position and evaluate a You have been advised that the Company has a marginal tax rate of 17% and all depreciable assets are depreciated using the straight-line depreciation method with zero salvage value. You have also been advised that the Companys target debt-to-equity ratio is 0.50. The Companys stock is trading at a beta of 0.75 and the yield-to-maturity on its bonds is 0. The risk-free rate is 6% and the expected market premium is 8%

(a) Calculate the cost of equity using the Capital Asset Pricing Model (“CAPM”) and Dividend Growth Model (“DGM”).

(b) Calculate the weighted average cost of capital for the Company. Assume the cost of equity to be the simple average of CAPM and DGM.
Answer #1

Solution:

Part a of the solution:

Calculation of Cost of equity as per CAPM:

As per the Capital Asset Pricing model (CAPM), Cost of equity is calculated using the following formula:

Cost of equity = (RE) = RF + ?* ( RM – RF )

Where RF = Risk free rate

? = Beta of the project

RM = Expected market premium

As per the data given in the question RF = 0.06 ; ? = 0.75 ;

Market risk premium = (Risk free rate – Return on market)= (RM - RF )= 0.08

Applying the same in the formula above we have

Required rate of return = (RE) = 0.06 + 0.75 * (0.08)

                                                           = 0.06 + 0.06           

                                                           = 0.06 + 0.06= 0.12

Thus the Cost of equity of the company as per CAPM = 0.12 = 12 %.

Calculation of Cost of equity as per Dividend Growth model:

As per the Dividend growth model value of a stock/share = D0 * (1+g) / (Ke – g)

Where D0 = Dividend paid in year 0 or last dividend paid

g = Growth rate

Ke = Cost of Equity

As per the Information given in the question:

D0 = $ 0.30; g= 5% ; Value of stock = $ 4.70 per share

Applying the above values in the formula for value of a stock as per the Dividend Growth model we have:

$ 4.70 = $0.30 ( 1 + 0.05 ) / (Ke – 0.05)

$ 4.70 = $0.30 ( 1.05 ) / (Ke – 0.05)

$ 4.70 = $0.315 / (Ke – 0.05)

$ 4.70 = $0.315 Ke – $ 0.01575

$ 0.315 Ke = $ 4.70 + $ 0.01575 = $4.71575

Ke = $ 4.71575 / $ 0.315 = 14.97 %

This Cost of Equity as per Dividend growth model = 14.97 %.

Part b of the Solution:

As required, assuming that the cost of equity is the simple average of CAPM & DGM

Cost of equity as per CAPM = 12 %

Cost of equity as per DGM = 14.97 %

Thus average cost of equity as per both the methods = ( 12 % + 14.97 %) / 2 = 26.97/2 = 13.485 %

Weighted average cost of capital is given by the following equation:

WACC = ( We * Ke ) + ( Wd * Kd [1-t])

Where We = Weight of equity ; Ke = Cost of equity    ; Wd = Weight of debt

Kd = Cost of debt ; t = tax rate ; WACC : Weighted average cost of capital

In the given question

Ke = 13.485 %    ; Kd =6 % ; Wd= 0.5   We = 0.5 ; t = tax rate = 0.17 = 17 % ;

Applying the above values in the formula we have

WACC = [0.5 * 13.485 % ]+[ 0.5 * ( 6 % ( 1 – 0.17 ) )]

            = 6.7425 % + [ 0.5 * 4.98 %]

            = 6.7425 % + 2.49 %

            = 9.2325 %

Thus the weighted average cost of capital of RonSIM Industries Limited is 9.2325 % = 9.23 % ( when rounded off to two decimal places)